OpenStudy (anonymous):

Question 7 Which of the following represents eighth root of x cubed in exponential form? x to the 8 thirds power x to the 3 eighths power 3x8 8x3 Question 8 Simplify (12 − 6i) − (−3 − 8i) 15 + 2i 15 − 14i 13 17 Question 9 Given the functions j(x) = x2 - 9 and k(x) = -x + 7, which operation results in a 3rd degree polynomial? Addition Subtraction Multiplication No operations

OpenStudy (anonymous):

@math&ing001 lol

OpenStudy (math&ing001):

\[\sqrt[8]{x ^{3}}=x ^{\frac{ 3 }{ 8 }}\] There are some few simple rules about exponential you should know about: \[x ^{-a}=\frac{ 1 }{ x ^{a} }\] \[\sqrt[a]{x}=x^{1/a}\]

OpenStudy (anonymous):

8?

OpenStudy (math&ing001):

You want me to guide you through ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

& was the last ones answer x^3/8 or x^8/3 ?

OpenStudy (math&ing001):

Use the second rule to find out ;)

OpenStudy (anonymous):

ok! so its 3/8 ? am i righttttt?

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

8?

OpenStudy (math&ing001):

For 8, first distribute (aka remove all brackets) then regroup terms alike

OpenStudy (anonymous):

ok! then ?

OpenStudy (math&ing001):

Then add those terms to each other.

OpenStudy (anonymous):

ok so is it 15+2i ?

OpenStudy (math&ing001):

Yeah ! You're a fast learner :D

OpenStudy (anonymous):

lol! 9?

OpenStudy (math&ing001):

J is second degree and k is first degree, you'll need to multiply them to get a higher degree.

OpenStudy (anonymous):

im confused..

OpenStudy (math&ing001):

You what the degree of a polynomial is ?

OpenStudy (anonymous):

no.

OpenStudy (math&ing001):

Suppose \[P(x) = x ^{6} + 5x ^{4} -10x ^{7} + 2x + 1\] Its degrree would be the highest power of all terms of the polynomial which is 7.

OpenStudy (anonymous):

can you just work this problem out for me? lol

OpenStudy (math&ing001):

It's against the Code of Conduct to just show answers http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i mean i want you to show me how to work it out then show me the correct answer?

OpenStudy (anonymous):

would it be addition?

OpenStudy (math&ing001):

Well, as I said before, you need to multiply j(x) and k(x) to get a 3rd degree polynomial, and that's basically the answer lol